# Asymptotic expansions for solutions of differential equations having a coalescing turning point and double pole, with an application to Legendre functions

### v3

@@ 5.1 Numerical Example
Because $\mathsf{P}^{-\mu}_{\nu}(t)$ has zeros on $[0,1)$, an envelope function is introduced to assess the relative error in this interval. Following [5, §6], therefore define
@@
Similarly, in Fig. 5 we plot $\Omega_{4}(\nu,\mu,t)$ for $\nu=50$ and $\mu=u\alpha$ with $\alpha=0.9$. Here, the turning point occurs at $\sigma=\sqrt{19}/10=0.43589\cdots$, which lies farther from the singularity. The accuracy throughout the interval remains good, though it is lower than in the case where the turning point is closer to the singularity. Compared to the previous parameter values, the turning points at $t=\pm\sigma$ are closer to each other. Since the present approximation is not uniform in a neighborhood of $t=-\sigma$, this increased proximity explains the reduction in overall accuracy. In such cases, a uniform expansion based on the new theory of coalescing turning points can yield a sharper approximation throughout the interval without increasing the number of terms; see [5, §5].
@@ Acknowledgment
Financial support from Ministerio de Ciencia e Innovación project PID2021-127252NB-I00 (MCIN/AEI/10.13039/ 501100011033/FEDER, UE) is acknowledged.

### v4

@@ 5.1 Numerical Example
Because $\mathsf{P}^{-\mu}_{\nu}(t)$ has zeros on $[0,1)$, an envelope function is introduced to assess the relative error in this interval. Following [Dunster:2025:LCT, §6], therefore define
@@
Similarly, in Fig. 5 we plot $\Omega_{4}(\nu,\mu,t)$ for $\nu=50$ and $\mu=u\alpha$ with $\alpha=0.9$. Here, the turning point occurs at $\sigma=\sqrt{19}/10=0.43589\cdots$, which lies farther from the singularity. The accuracy throughout the interval remains good, though it is lower than in the case where the turning point is closer to the singularity. Compared to the previous parameter values, the turning points at $t=\pm\sigma$ are closer to each other. Since the present approximation is not uniform in a neighborhood of $t=-\sigma$, this increased proximity explains the reduction in overall accuracy. In such cases, a uniform expansion based on the new theory of coalescing turning points can yield a sharper approximation throughout the interval without increasing the number of terms; see [Dunster:2025:LCT, §5].
@@ Acknowledgment
Financial support from Ministerio de Ciencia e Innovación project PID2024-159583NB-I00 (MICIU/ AEI / 10.13039/501100011033 / FEDER, UE) is acknowledged.